### v1

# Wave packets, “negative times” and the elephant in the room

Abstract
Controversy surrounding the “tunnelling time problem” stems from the seeming inability of quantum mechanics to provide, in the usual way, a definition of the duration a particle is supposed to spend in a given region of space. For this reason, the problem is often approached from an “operational” angle. One such approach uses the position of the transmitted wave packet in order to infer the duration the particle spends in the barrier. Here we replace the barrier with a tuneable Mach-Zehnder interferometer (MZI). With this analogy one is able, at least in principle, to achieve any advance or delay of the wave packet sent to the chosen outgoing port. The Uncertainty Principle prevents one from combining the durations spent in each arm the MZI into a meaningful duration when both arms are engaged. There is no justification for invoking “superluminal” or “negative” times, since the particle is able to arrive at the same position (and with a higher probability) if the same initial state propagates through only one arm of the MZI. The same is true, we argue, in the case of tunnelling, where the transmitted wave packet results from destructive interference between multiple copies of the free state, delayed relative to the free propagation.
I Introduction
It is a commonplace to state that the tunnelling time problem, i.e., the problem of determining the duration ${\tau}$ a tunnelling particle spends in the barrier, continues to cause controversies as of now. In standard quantum mechanics, there have been at least three main approaches, all trying to extend classical reasoning to an essentially quantum context. In 1932 McColl proposed deducing the duration from the position of the transmitted wave packet WP0 . The method was later extended in WP1 -WP2 . Thirty five years later Baz’ L1 proposed equipping the scattered particle with a magnetic moment (spin), and determining ${\tau}$ by dividing its angle of rotation in a small magnetic field by the Larmor frequency $\omega_{L}$. Finally, in 1982 Büttiker and Landauer suggested extracting the duration spent in the barrier from the adiabatic limit reached in a small oscillatory field added to the barrier BL . The literature on the subject is extensive, and we refer the reader to various reviews (see, e.g., REV1 -REV7 ). None of the three methods mentioned so far are free from problems (for a recent discussion of the Büttiker-Landauer time see DSBL ) Below we will be interested mostly in the wave packet technique WP0 -WP2 , to which we will return after a brief digression.
II The elephant in the room
Perhaps the most straightforward analysis the problem can be given in the case of the Larmor (Baz’) clock DSL1 . Feynman paths connecting the particle’s initial and final states can be grouped into “pathways” according to the duration ${\tau}$ a path spends on the region $\Omega$ where the magnetic field is introduced. By summing the corresponding path amplitudes one obtains the amplitude $\mathcal{A}({\tau})$ of spending in $\Omega$ $\tau$ seconds, The spin’s final state is, therefore, a superposition of the copies of its initial state, rotated by by all $\omega_{L}{\tau}$.
One basic difficulty is now apparent. Tunnelling is an interference phenomenon. The small tunnelling amplitude is a result of destructive interference between the pathways. The Uncertainty Principle FeynL - the elephant in the room where a tunnelling time discussion is held- forbids knowing the path taken by the particle without destroying the interference pattern, i.e., by destroying the very tunnelling event one wishes to study.
The difficulty is already evident if the number of pathways reduced to two DSL2 . Consider a Mach-Zehnder interferometer (MZI), with a small magnetic field introduced in such a way that a particle with a spin spends in it durations ${\tau}_{1}$ and ${\tau}_{2}$, if travelling via its first and the second arm, respectively. On exit from the interferometer, at a location which the particle can reach via each arm with amplitudes $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, the spin’s angle of rotation is given by DSL2
$\varphi=\omega_{L}\text{Re}\left[\frac{{\tau}_{1}\mathcal{A}_{1}+{\tau}_{2}\mathcal{A}_{2}}{\mathcal{A}_{1}+\mathcal{A}_{2}}\right].$ (1)
With only a few apriori restrictions on the amplitudes $\mathcal{A}_{1,2}$, the real part of the “complex time” in the square brackets is not an acceptable duration. It can be zero, prompting suspicion of violation of Einstein’e relativity, or negative, or even exceed the time it takes the particle to travel to the chosen location DSL2 , Whoever wishes to avoid unsavoury sensationalism, is forced to admit the failure of inferring a meaningful duration by combining ${\tau}_{1}$ and ${\tau}_{2}$, unless the interference between the paths is destroyed. This must be, we note, a consequence of the Uncertainty Principle FeynL .
Below we will show that the same problem will occur if, instead of employing Larmor precession, one chooses to monitor the position of the wave packet exiting the interferometer. As with the Larmor clock, the Uncertainty Principle is resonsible- the same elephant in a different room.
III Wave packet delays
Consider a Gaussian wave packet (WP)
$G(x)=(\pi\Delta x^{2}/2)^{-1/4}\exp(-x^{2}/\Delta x^{2})$ (2)
injected in into the Mach-Zehnder interferometer shown in Fig.1a.
Figure 1: a) A wave packet is divided into two parts which are recombined agaim while passing through a pair of beam splitters(BS). A delay by ${\tau}$ sec. occurs for the part passing through the right route. A wave packet travelling to detector $D1$ consists of two parts (cf. Eq.(3) which may or may not overlap, depending on $v$, ${\tau}$, and the initial wave packet’s width $\Delta x$. Their sum can result in a much smaller advanced Gaussian. b) A path diagram of the system in a) and the four amplitudes $\mathcal{A}_{i}$. The paths leading to different final states $D_{1}$ and $D_{2}$ never interfere. c) An equivalent quantum measurement problem (see Sect.IV.
The WP, propagating at velocity $v$, splits at the first beamsplitter ($BS$), experiences a delay by $\tau$ secs in the MZI’s right arm, its parts recombined after the second $BS$. For simplicity, we neglect WP’s spreading and set the new origin, $x=0$, at the centre of the WP, propagating along the left arm, (cf. Fig.1a). The part of the wave function travelling towards detector $D_{1}$, is given by
$G_{1}(x)=\mathcal{A}_{1}G(x)+\mathcal{A}_{2}G(x+v{\tau}),$ (3)
where $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are the path amplitudes, determined by the properties of the beamsplitters, etc. The width of the initial state now controls the amount of interference between the two terms in (3).
[…]
In the opposite limit $\Delta x>>v{\tau}$ one finds $G(x+v{\tau})\approx G(x)$, interference between the pathways is duly restored and in, accordance with the Uncertainty Principle, the probability to reach $D_{1}$ is $P\approx|A_{1}+A_{2}|^{2}$. There is not a single $x$, such that the particle found there can be said to travel one path, and not another. Moreover, its is a well known property of Gaussians Vaid that as $\Delta x\to\infty$,
$P(x)\xrightarrow[\Delta f_{2}\to 0]{}|\mathcal{A}_{1}+\mathcal{A}_{2}|^{2}\exp\left[-2\frac{(x-\overline{x})^{2}}{\Delta x^{2}}\right],$ (4)
with the positions of the peak (and the centre of mass) of the Gaussian in the r.h.s. of (4) given by an expression somewhat similar to (1)
$\overline{x}\equiv-\text{Re}\left[\frac{v{\tau}\mathcal{A}_{2}}{\mathcal{A}_{1}+\mathcal{A}_{2}}\right].$ (5)
Convergence of $P(x)$ to the single-peak form (4), described by the Catastrophe Theory CAT , can be shown to be point-wise DSA .
Next we note that one can, at least in principle, choose the delay $\tau$ and the amplitudes $\mathcal{A}_{1,2}$ so as to give the distance between the peak propagating via both arms, and the peak travelling via the left arm only, $\overline{x}$, any desired value.
IV Any delay is possible
The system shown in Figs.1a and 1b is described by the total of four paths, two for each final state $D_{1}$ and $D_{2}$. The four complex valued path amplitudes, $\mathcal{A}_{i}$, $i=1,2,3,4$ need to satisfy the probability conservation rules.
$\sum_{i=1}^{4}|\mathcal{A}_{i}|^{2}=|\mathcal{A}_{1}+\mathcal{A}_{2}|^{2}+|\mathcal{A}_{3}+\mathcal{A}_{4}|^{2}=1,$ (6)
However, little restriction is imposed on the signs of their real and imaginary parts. Thus, for a given real $y$, one can always find $\mathcal{A}_{i}$ such that
$\text{Re}\left[\frac{y\mathcal{A}_{1}}{\mathcal{A}_{1}+\mathcal{A}_{2}}\right]=z$ (7)
for any desired $z$. Indeed, Eq.(3) also describes the state of a Gaussian von Neumann pointer vN , set up to measure a projector, $\hat{B}=y|b_{2}{\rangle}{\langle}b_{2}|$, $y=-v{\tau}$, for a two-level system (qubit), pre- and post-selected in some states $|I{\rangle}$ and $|D_{1}{\rangle}$, respectively DSS . In this case, the amplitudes of the four paths shown in Figs. 1b and 1c, are known to be given by DSS (neither the pointer, nor the qubit, have own dynamics)
$\mathcal{A}_{1}={\langle}D_{1}|b_{1}{\rangle}{\langle}b_{1}|I{\rangle},$ (8)
$\mathcal{A}_{2}={\langle}D_{1}|b_{2}{\rangle}{\langle}b_{2}|I{\rangle},$
$\mathcal{A}_{3}={\langle}D_{2}|b_{1}{\rangle}{\langle}b_{1}|I{\rangle},$
$\mathcal{A}_{4}={\langle}D_{2}|b_{2}{\rangle}{\langle}b_{2}|I{\rangle}.$
Constructed in this way, the amplitudes (8) automatically satisfy the conditions (6), and one only needs to show that , for a given initial state $|I{\rangle}$, and a basis $|b_{i}{\rangle}$, it is always possible to find a pair of orthonormal final states $|D_{i}{\rangle}$. There are various choices, including the one which will keep $\mathcal{A}_{i}$ real, e.g.,
${\langle}b_{1}|D_{1}{\rangle}=N^{-1}/{\langle}b_{1}|I{\rangle}^{*},\quad\quad\quad\quad\quad\quad\quad$ (9)
${\langle}b_{2}|D_{1}{\rangle}=N^{-1}\left(\frac{z}{y-z}\right)/{\langle}b_{2}|I{\rangle}^{*},\quad\quad\quad$
${\langle}b_{1}|D_{2}{\rangle}={\langle}b_{2}|D_{1}{\rangle}^{*},\quad\quad\quad\quad\quad\quad\quad\quad\quad$
${\langle}b_{2}|D_{2}{\rangle}=-{\langle}b_{1}|D_{1}{\rangle}^{*},\quad\quad\quad\quad\quad\quad\quad\quad$
$N=\sqrt{\frac{1}{|{\langle}b_{1}|I{\rangle}|^{2}}+\frac{z^{2}}{(y-z)^{2}|{\langle}b_{2}|I{\rangle}|^{2}}}.\quad\quad$
In the special case $|I{\rangle}=(|b_{1}{\rangle}+|b_{2}{\rangle})/\sqrt{2}$ the four amplitudes in (8) become
$A_{1}=\frac{1}{\sqrt{2[1+\frac{z^{2}}{(y-z)^{2}}]}},\quad\quad\quad\quad\quad\quad$ (10)
$A_{2}=\frac{z}{y-z}A_{1},\quad\quad\quad\quad\quad\quad\quad\quad\quad$
$A_{3}=A_{2},\quad A_{4}=-A_{1}.\quad\quad\quad\quad\quad\quad$
For a particle in Fig.1a we put $y=-v\tau$, and, wishing only to illustrate a principle, assume that the amplitudes in Eqs (8) can somehow be realised in practice. We then note that for $z=0$. $A_{2}=A_{3}=0$, and the particle arrives in $|D_{1}{\rangle}$ via $|b_{1}{\rangle}$ (the first path in Fig.1b), and in $|D_{2}{\rangle}$ via $|b_{2}{\rangle}$ (the fourth path in Fig.1b). Similarly, for $z=-v\tau$, both final states $|D_{1,2}{\rangle}$ are reached via $|b_{2}{\rangle}$.
V Advancing the peak
Now the following experiment is, at least in principle, possible. First one removes the beamsplitters in Fig.1a, let a broad Gaussian WP travel towards $D_{1}$ along the left arm of the MZI until time $T$, and determines the position of its peak (or of the centre of mass). Next the beamsplitters are re-introduced, so a part of the initial WP is diverted in through the right arm, where it is delayed by ${\tau}$, and then recombined with the part travelling the right arm. The result is the Gaussian in Eq.(4), whose peak’s position $\overline{x}$ is then compared to that of the WP travelling along the left arm only, $|G(x)|^{2}$. As was demonstrated above, with an appropriate choice of the amplitudes $\mathcal{A}_{1.2}$ the peak can be made to lie ahead of the free one, $\overline{x}>0$, (or arrive earlier at a fixed detector, if one so wishes). To observe this “speed up” produced by the delay in the right arm, one does not even need to make initial WP very broad. Transformation of the of a twin-peak density $P(x)$ into a single-peak form as the width $\Delta x$ increases is followed in Fig.2, with a reasonable approximation to Eq.(4) is obtained already for $\Delta x/v{\tau}=5$, as is shown in Fig.3.
Figure 2: Contour plot of the probability density $P(x)=|\mathcal{A}_{1}G(x)+\mathcal{A}_{2}G(x+v{\tau})|^{2}$ vs. $x$ and $\Delta x$ for the case where the second WP is delayed, for $\mathcal{A}_{1}=0.5883$ and $\mathcal{A}_{2}=-0.3922$. As the width $\Delta x$ increases, the delayed maximum initially at $x=-v{\tau}$ disappears, accompanied by the minimum, located for $\Delta x<<v{\tau}$ at $x=-v|{\tau}|/2$. The surviving maximum is pushed forward, and asymptotically approaches $\overline{x}=2$ [cf. (5)], indicated by the dotted line. Also shown is the centre of mass (COM) of the density, which coincides with the density’s peak only in the limit $\Delta x\to\infty$,
Figure 3: Comparison between the normalised to unity density $|G_{1}(x)|$ in Eq.(3) and that obtained from Eq.(4) with $\overline{x}$ given by Eq.(17) for $\Delta x/v{\tau}=5$. The amplitudes $\mathcal{A}_{1,2}$ are the same as in Fig.2, and the advancement of the peak of about $1.35v{\tau}$ is somewhat smaller than $2v{\tau}$, achieved for $\Delta x/v{\tau}\to\infty$.
Furthermore, for $\overline{x}>0$ the second of Eqs.(10) yields
$\frac{\mathcal{A}_{2}}{\mathcal{A}_{1}}=-\frac{\overline{x}}{\overline{x}+v{\tau}}<0,$ (11)
so one converts a delay into advancement by means of destructive interference. In particular, as $\overline{x}\to\infty$, $\mathcal{A}_{2}\to-\mathcal{A}_{1}$, which makes the Gaussian (4) very small, and detection by $D_{1}$ highly unlikely.
For the ratio between $|G(x)|^{2}$ and that in Eq.(4) one finds
$\frac{P(x)}{|G(x)|^{2}}=|\mathcal{A}_{1}|^{2}\left|1-\frac{\overline{x}}{\overline{x}+v{\tau}}\frac{G(x+vt)}{G(x)}\right|^{2}$ (12)
where we have used (11). For any $x>0$ and $\overline{x}>0$ the ratio (12) is less than unity. In other words, the density (4), if advanced, will always fit under the front tail of the density, transmitted to $D_{1}$ through the left arm only (see Fig.4).
Figure 4: Comparison between the density $|G_{1}(x)|$ in Eq.(3) and the one obtained for the initial WP travelling along the left arm only. The parameters are the same as in Figs. 2 and 3.
Thus, for $\Delta x>>vt$ most particles are directed towards the detector $D_{2}$, where interference between the paths is constructive. The probability density is still given by Eqs.(4)-(5), with the amplitudes $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ replaced with $\mathcal{A}_{3}$ and $\mathcal{A}_{4}$, respectively. The peak of the WP lies a distance at $vt/2$ behind the peak propagating through the left arm only, and the probability of detection by $D_{2}$ tends to 111.
VI Delaying the peak
One can also prepare a setup where the transmitted peak lags behind the peak of the WP travelling through the left arm, $\overline{x}<0$ , by choosing
$\frac{\mathcal{A}_{2}}{\mathcal{A}_{1}}=-\frac{|\overline{x}|}{|\overline{x}|-v{\tau}}.$ (13)
The ratio is positive, and interference is constructive, for as long as the peak still lies ahead of the density $|G(x+v{\tau})|^{2}$, corresponding to the right arm only, i.e., for $|\overline{x}|<v{\tau}$. If one tries to delay by more than $v{\tau}$ interference becomes destructive again. It is easy to see that the peak, delayed in this manner, fits under the rear tail of the delayed Gaussian , since
$\frac{P(x)}{|G(x+v{\tau})|^{2}}=|\mathcal{A}_{2}|^{2}\left|1-\frac{|\overline{x}|-v{\tau}}{|\overline{x}|}\frac{G(x)}{G(x+vt)}\right|^{2}<1\quad\quad$ (14)
for all $x<-v{\tau}$. As with the advancement, for a large delay, $|\overline{x}|>>vt$, one has $\mathcal{A}_{2}\to-\mathcal{A}_{1}$, and detection by $D_{1}$ becomes highly improbable.
VII “Superluminality”, “negative times” and “abnormal delays”
So far there was nothing unusual in our analysis of the MZI interference. In neither case the particle was found in a place where it would not arrive if the same initial state were propagated along only one arm of the interferometer.
Spurious sensationalism will enter the discussion if, following WP0 , one decides to convert spatial displacements into temporal durations. But, might object a fastidious reader, surely the small peak, advanced by $\overline{x}$, would arrive at a fixed detector $\overline{x}/v$ seconds earlier than the peak of the WP travelling along the left arm only? Is it not an observable time interval? Yes, but one cannot deduce from this temporal delay the duration the particle spends inside the MZI. the duration the particle spends the MZI.
This is a kind of reasoning which, in our opinion, one should try to avoid. Let the distance between the beamsplitters measured along the left arm of the MZI be $L$. Then the maximum of the wave packet, travelling only along the left arm, will cover it in $L/v$ seconds. If the peak (that is the maximum) of the wave packet transmitted across both arms and exiting along the destructive interference port is found a distance $\overline{x}$ ahead, the particle “must have spent between the BSs a shorter time”,
${\tau}_{inside}=L/v-\overline{x}/v.$ (15)
But one can choose the delay ${\tau}$, $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, in order to ensure that the transmitted peak, resulting from the interference between the forward tails of the Gaussians passing through the left and right arms, leaves the MZI just as the maximum of the initial Gaussian enters it, $\overline{x}=L$. Then one may be tempted to claim that the MZI has been crossed infinitely fast, in defiance of Einstein’s relativity. This, we argue, is not true, since if we block the right arm of the MZI, a wavepacket traveling through the left arm only would also lead to the detection of a particle at the position where the peak of the interfering wavepacket appears, and would do so with a higher probability [cf. Eq.(12)]. initial state were propagated along the left arm of the MZI at a perfectly subluminal speed $v$.
Worse still, arranging for $\overline{x}>L$ would make the duration ${\tau}_{inside}$ negative, so the transmitted peak leaves the MZI before the incident one enters it. Such a behaviour was reported, for example in BEF . Again, the same argument suggests that that nothing untoward happens also in this case. Besides, the notion of a “negative duration” is not defined, just as the concept of “negative probability”, which cannot be related to frequencies FeynNEG . Rather than claiming “experimental evidence of quantum negative time”, one should perhaps question the method leading to this bizarre notion.
Finally, the same applies to abnormally long delays, $\overline{x}<<-v{\tau}$, ${\tau}_{inside}>>L/v$. One would also detect such particles in the tail of a broad wave packet, propagating only via the right arm of the MZI, where it experiences a much shorter delay $\tau$.
VIII Conclusions
It is probably fair to say that few authors authors, with possible exception of Nimtz and Stahlhofen NS , are prepared to accept that classically forbidden quantum events, such as tunnelling, fall outside the remit of special relativity. More often the apparent speed up of the transmitted particle is related to some form of reshaping, which destroys the causal relationship between the incoming and transmitted peaks of the wave packets involved REV1 ,BW . Naive, and purely classical examples of reshaping include the train analogies, discussed in WIN , and the paper-and-scissors metaphor of DSE .
A more realistic mechanism, considered here, relies on interference between the parts of a broad wave packet travelling through different arms of a two-way interferometer. By arranging the path amplitudes, one can, in principle, place the transmitted peak anywhere before, or behind, the peak of the WP travelling through one arm only. With a delay of ${\tau}$ secs experienced in one arm, overall advancement of the transmitted peak requires the interference to be destructive. When recombined, the parts of the WP have opposite signs, and cancel each other, leaving only a small advanced peak. It is a peculiar property of the Gaussians that a broad advanced WP, although greatly reduced, repeats the shape of the incident pulse. The same mechanism, we note, is responsible for delaying the transmitted WP by more that ${\tau}$.
Spurious controversies may arise if one uses the position of the transmitted peak hoping to deduce the duration spent by the particle inside the interferometer. It is, however, obvious there is neither “superluminal behaviour”, or the possibility of “time travel” where ${\tau}_{inside}$ in Eq.(15) turns negative. The simple reason is that that the particle is perfectly able to arrive at the same time in the same place in a perfectly pedestrian manner, i.e., if the same initial state propagates through just one arm, at a purely subliminal velocity $v$. It is comforting, although not necessary for our argument, to note that the corresponding probability is reduced, whenever both arms of the interferometer are engaged, and the rest of the particles are re-directed towards the other detector $D_{2}$.
We note further that at the centre of controversy is quantum interference or, more precisely, the Uncertainty Principle. The Principle forbids revealing the path travelled by a particle without destroying interference between the paths in the process FeynL . Similarly, with the interference intact, it does not allow one to combine the durations spent in each arm into a single meaningful duration, describing the passage when both arm are available to the particle.
This is, in a nutshell, the tunnelling time problem. It is easy to show DSE , NOT that in the case of a potential barrier the transmitted state is given by a similar superposition of various copies of the freely propagating wave packet, subjected to various spatial delays. Because of destructive interference, the tunnelling amplitude is small, although individual terms in the superposition are not. The two-way analogue studied here captures the main features of the phenomenon, and we expect most of the above conclusions to apply also in the case of tunnelling. It is difficult to imagine, given the simplicity of our example, how a further explanation can be given to such a basic destructive interference phenomenon.
Acknowledgements
DS acknowledges financial support by the Grant PID2021-126273NB-I00 funded by MICINN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”, as well as by the Basque Government Grant No. IT1470-22.